OpenStudy (anonymous):

solve equation -2tan(theta) = 2root3

OpenStudy (anonymous):

\[-2 \tan(\theta) = 2\sqrt3 \rightarrow \tan(\theta) = \frac{2\sqrt3 }{-2} \] \[\tan(\theta) =-\sqrt3 = -\tan \frac{\pi}{3} = \tan (\pi- \frac{\pi}{3} ) = \tan (\frac{2\pi}{3} )\] \[\tan(\theta) = \tan (\frac{2\pi}{3} ) \rightarrow \theta =\frac{2\pi}{3}\]